Sarah spends 1/4 of her income for transportation, 1/6 of her income for food, and 1/3 of her income for rent and utilities. What fractional part of Sarah's income is left for all of her other expenses?
Answer by stanbon(75887)

Sarah spends 1/4 of her income for transportation, 1/6 of her income for food, and 1/3 of her income for rent and utilities. What fractional part of Sarah's income is left for all of her other expenses?
---------
1-1/4-1/6-1/3
Lcd is 12
= 12/12 - 3/12 -2/12 -4/12
= (12-9)/12
= 3/12
= 1/4 is left for all her other expenses
=======================
Cheers,
Stan H.